# Debias Can be Unreliable: Mitigating Bias Issue in Evaluating Debiasing Recommendation

### v1

@@ Abstract
Recent work has improved recommendation models remarkably by equipping them with debiasing methods. Due to the unavailability of fully-exposed datasets, most existing approaches resort to randomly-exposed datasets as a proxy for evaluating debiased models, employing traditional evaluation scheme to represent the recommendation performance. However, in this study, we reveal that traditional evaluation scheme is not suitable for randomly-exposed datasets, leading to inconsistency between the Recall performance obtained using randomly-exposed datasets and that obtained using fully-exposed datasets. Such inconsistency indicates the potential unreliability of experiment conclusions on previous debiasing techniques and calls for unbiased Recall evaluation using randomly-exposed datasets. To bridge the gap, we propose the Unbiased Recall Evaluation (URE) scheme, which adjusts the utilization of randomly-exposed datasets to unbiasedly estimate the true Recall performance on fully-exposed datasets. We provide theoretical evidence to demonstrate the rationality of URE and perform extensive experiments on real-world datasets to validate its soundness.
1 Introduction
@@
Recommendation models suffer from various biases such as selection bias (Marlin et al. 2007; Hemminki and Sihvo 1993; Huang, Oosterhuis, and de Rijke 2022; Vernade and Cappé 2015) and popularity bias (Abdollahpouri and Mansoury 2020; Zhang et al. 2021; Zhu et al. 2021; Abdollahpouri et al. 2021), leading to deviations from true user preferences even harmful impacts to Web services (Chen et al. 2023; Klimashevskaia et al. 2023; Klimashevskaia 2022; Kowald, Schedl, and Lex 2020), such as amplifying the long-tail effect and hurting user trust (Abdollahpouri, Burke, and Mobasher 2017; Abdollahpouri et al. 2019; Abdollahpouri 2019). Undoubtedly, user’s true preferences can be accurately measured by the fully-exposed dataset which includes users’ feedback (e.g., ratings) across all items. The gold standard for evaluating debiasing techniques is thus utilizing such a fully-exposed dataset to assess the performance of recommendations (Gao et al. 2022a). However, due to the unavailability of fully-exposed datasets in practice (Gao et al. 2022b), randomly-exposed datasets, which contain users’ feedback on some randomly selected items, are widely employed for debiasing evaluation (Liu et al. 2020; Yuan et al. 2019; Rosenfeld, Mansour, and Yom-Tov 2017; Bonner and Vasile 2018; Schnabel et al. 2016; Wang et al. 2019; Ai et al. 2018; Chen et al. 2021). Notably, the amount of feedback in randomly-exposed datasets is also extremely limited, because randomly exposing items to collect user feedback is highly expensive, and the collection process can severely damage a platform’s revenue.
Figure 1: Preliminary analysis on KuaiRec, an only existing fully-exposed dataset, validating the inconsistency between Recall@$\overline{K}$ on the randomly-exposed data ($l$ = 80 samples per user) and Recall@$K$ on the fully-exposed data ($L$ = 3327 samples per user) when $K$ is small. We rank four AutoDebias models with different hyperparameters w.r.t. Recall@$\overline{5}$ and Recall@$K$, respectively. Notably, $\overline{K}\in$ [1,2,…$l$], $K\in$ [1,2,…$L$], where $l<<L$.
Previous work calculate widely-used Recall metric on randomly-exposed datasets in the same way as on fully-exposed datasets, which we refer to as the traditional evaluation scheme in subsequent descriptions. Nevertheless, we claim that traditional evaluation scheme is not suitable for randomly-exposed datasets when evaluating debiasing recommendation. Specifically, leveraging the only existing full-exposed dataset, KuaiRec (Gao et al. 2022b), our preliminary analysis reveals significant disparities between the evaluation results obtained from the randomly-exposed dataset and those derived from the fully-exposed dataset. Figure 1 is a quick example that presents the rankings of different models (i.e., AutoDebias models with different hyperparameters) based on the Recall@$\overline{5}$ on the randomly-exposed dataset and Recall@$K$ on the fully-exposed dataset. As can be seen, the model ranking deduced from Recall@$\overline{5}$ is different from those derived from Recall@$K$ when $K\leq 100$. Given that user typically browse only a limited number of items during an online deployment stage, Recall@$K$ with a small value of $K$ (e.g., $K=50$) on the fully-exposed dataset represents desired debiasing quality and correct model ranking. Consequently, the conclusion derived from the traditional evaluation scheme on the randomly-exposed dataset such as Recall@$\overline{5}$ might be unconvincing.
@@
To comprehensively investigate the underlying factors contributing to the inconsistency between the Recall@$\overline{K}$ on the randomly-exposed dataset and Recall@$K$ on the fully-exposed dataset, we undertake a theoretical analysis (cf. Theorem 3) of the evaluation performed on a randomly-exposed dataset. Our analysis demonstrates that the Recall@$\overline{K}$ on the randomly-exposed dataset (e.g., Recall@$\overline{5}$) has a weak correlation with the Recall@$K$ using a small value of $K$ on the fully-exposed dataset (e.g., Recall@50). Although the correlation between Recall@$\overline{K}$ and Recall@$K$ could become stronger as the value of $K$ increases, the performance cutting at a small value of $K$ is more important in practice (Chen et al. 2023). In addition, we conduct extensive experiments on real-world datasets to further validate the variance of such correlation (cf. Figure 1).
Given that the traditional evaluation scheme on the randomly-exposed dataset has limited capacity to reflect the desired Recall@$K$ performance on the fully-exposed dataset, it is inadequate to conclude whether previous debiasing methods have effectively mitigated biases. The observed performance gain of debiasing methods on the traditional evaluation scheme can be attributed to optimizing Recall@$K$ under a large value of $K$ on the fully-exposed dataset using techniques that lack genuine debiasing effects. For instance, hard negative sampling that theoretically optimizes Recall@$K$ with a large $K$ on fully-exposed datasets (Shi et al. 2023), can outperform the representative debiasing method (Chen et al. 2021) under the traditional evaluation scheme (see Table 3). Considering the cost of constructing fully-exposed dataset, unbiasedly estimating the Recall@$K$ performance with a small value of $K$ on the fully-exposed dataset from randomly-exposed datasets is crucial for the evaluation of debiasing methods.
Figure 2: Illustration of URE. In randomly-exposed data, $n$ is the number of positive items with existing feedback and $m$ is the number of positive items with existing feedback ranked within top-$K$ of all items.
@@
Towards this goal, we propose the Unbiased Recall Evaluation (URE) scheme as illustrated in Figure 2. We set the target as estimating the Recall@$K$ on fully-exposed data, i.e., the ratio of items with positive feedback ranked before position $K$, denoted as $\frac{M}{N}$. Considering that a randomly-exposed dataset only includes user feedback on randomly selected items, we can treat it as a representative subset of the fully-exposed dataset, wherein both positive and negative feedback are missing at random. In URE, we sort the model’s prediction scores of all candidate items and utilize the score of $(K+1)$-th item as the threshold to calculate the positive ratio on the randomly-exposed dataset, where the positive ratio is the proportion of items with positive feedback in the randomly-exposed dataset that rank before the threshold. Hence such a positive ratio can be used to unbiasedly estimate $\frac{M}{N}$ on fully-exposed dataset. This is due to the models’ prediction information of all candidate items, which is not considered by traditional evaluation scheme. Through theoretical analysis and extensive experiments, we validate that URE can unbiasedly estimate the Recall@$K$ on fully-exposed data with a desired value of $K$.
In this work, our contributions are threefold.
@@
  • We theoretically and empirically analyze the relation between the Recall@$\overline{K}$ on the randomly-exposed dataset and the Recall@$K$ on the fully-exposed dataset.
  • We propose the URE scheme to unbiasedly estimate the target Recall@$K$ performance with randomly-exposed data, and conduct extensive theoretical and empirical analysis, validating the effectiveness of the URE scheme.
2 Preliminary
Table 1: Notations and definitions.
Notations Annotations
$D_{full}$ A fully-exposed dataset
$D_{rand}$ A randomly-exposed dataset
$L/l$ The amount of feedback from a single user
$K$ Top-K on $D_{full}$, $K\in$ [0,1,…$L$]
$\overline{K}$ Top-K on $D_{rand}$, $\overline{K}\in$ [0,1,…$l$] ($l<<L$)
Recall@$K$ Recall metric at position $K$ on $D_{full}$
Recall@$\overline{K}$ Recall metric at position $\overline{K}$ on $D_{rand}$
$\widehat{\text{Recall@}K}$ The estimation of $\text{Recall}@K$ by URE scheme
Given a user ($u$) set and an item ($i$) set, fully-exposed dataset $D_{full}$ includes the user feedback $r_{u,i}$ (e.g., ratings and clicks) on all candidate items, which is usually unavailable in practice. The only existing $D_{full}$ for research is KuaiRec (Gao et al. 2022b). All candidate items can be categorized into items with positive feedback that user likes ($r_{u,i}=1$) and items with negative feedback ($r_{u,i}=0$) based on user’s true interests. For the evaluation of debiasing recommendation, researchers mostly resort to randomly-exposed datasets $D_{rand}$ as a proxy to evaluate debiasing methods. However, randomly-exposed dataset $D_{rand}$ collects the user feedback on some randomly selected items ($\in$ all candidate items) through a costly randomized intervention experiment (Liu et al. 2020; Yuan et al. 2019; Rosenfeld, Mansour, and Yom-Tov 2017; Bonner and Vasile 2018; Schnabel et al. 2016; Wang et al. 2019; Ai et al. 2018; Chen et al. 2021), i.e., a large amount of user feedback is ramdomly missing. During the evaluation stage, recommendation models ($\hat{y}=M(u,i)$) rank the items with feedback based on $\hat{y}$ to generate a personalized ranking list $Rank_{u}$ for each user. In traditional evaluation scheme on $D_{full}$, we can obtain the true performance of debiased recommendation models by calculating the Recall@$K$. Notably, The Recall metric is the most important metric in the candidate generation stage of industrial recommendation systems. It is also a commonly used metric in research on debiasing recommendations. Recall@$K$ on $D_{full}$ is obtained by the equation 1.
$\text{Recall}@K=\frac{|{k\in Rank_{u}+:k\leq K}|}{|Rank_{u}+|},K\in[1,2,...,L],$ (1)
where the $Rank_{u}+$ collects the positions of all items with positive feedback within $Rank_{u}$.
Unfortunately, as mentioned above, fully-exposed dataset is unavailable on most recommendation platforms. The Recall@$K$ in the equation 1 is not supported if the feedback of any candidate item is missing. For the evaluation of debiasing recommendation, researchers employ traditional evaluation scheme on $D_{rand}$ to calculate the Recall performance, denoted as Recall@$\overline{K}$.
Figure 3: The Pearson correlation coefficients or Kendall coefficients between Recall@$\overline{K}$ on $D_{rand}$ and Recall@$K$ on $D_{full}$. We highlight the value of $K$ when each curve reaches its maximum correlation coefficient as $K_{max}$. (a) The effect of $\overline{N}$, where $\overline{K}$ is fixed to 5. (b) The effect of $\overline{K}$, where $\overline{N}$ is fixed to 80. (c) The effect of $\overline{N}$, where $\overline{K}$ is fixed to 5. The correlations are statistically significant with $p$ $<$ 0.01.
$\begin{split}\text{Recall}@\overline{K}=\frac{|{k\in\overline{Rank_{u}}+:k\leq\overline{K}}|}{|\overline{Rank_{u}}+|},\\ \overline{K}\in[1,2,...,l],\quad l<<L,\end{split}$ (2)
where $\overline{Rank_{u}}+$ contains the positions of items with positive feedback of user $u$ in $\overline{Rank_{u}}$. However, in this work, we find that Recall@$\overline{K}$ on the randomly-exposed dataset is inconsistent with Recall@$K$ on fully-exposed dataset. Therefore traditional evaluation scheme is not suitable for randomly-exposed datasets. We will detail the inconsistency and unreliability of such a traditional evaluation scheme in the following sections.
@@ 2 Preliminary
In this section, we first conduct theoretical analysis to discover why the traditional evaluation scheme on randomly-exposed data is unreliable by inspecting the relationship between Recall@$K$ metric on fully-exposed data and Recall@$\overline{K}$ metric on randomly-exposed data. Then, we substantiate their correlation through empirical validation.
3.1 Theoretical Guarantee
In this subsection, we give the theoretical derivation of the relationship between the Recall@$K$ on $D_{full}$ and the Recall@$\overline{K}$ on $D_{rand}$.
@@ 3.1 Theoretical Guarantee
$E_{\mathcal{R}}\left[E_{\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}]-\text{Recall@}K\right]=0,$ (3)
@@
The proof of Theorem 3 can be found in Appendix A.2. From Theorem 3, we can observe that the Recall@$\overline{K}$ on $D_{rand}$ will exhibit a strong correlation with the Recall@$K$ on $D_{full}$ if the value of $K$ is around $\frac{{N}}{\overline{N}}\cdot\overline{K}$. It should be noted that $\overline{N}$ is much smaller than $N$ since the randomly-exposed dataset is a sparsely sampled subset of the fully-exposed dataset. Therefore, given a specific value of $\overline{K}$, the Recall@$\overline{K}$ on $D_{rand}$ only exhibits a strong correlation with the Recall@$K$ with much larger $K$ values on $D_{full}$. However, given that user typically browse only a limited number of items during an online deployment stage, Recall@$K$ with a small value of $K$ (e.g., $K=50$) on the $D_{full}$ represents desired debiasing quality and correct model ranking. Although the correlation between Recall@$\overline{K}$ and Recall@$K$ could become stronger as the value of $K$ increases, the performance cutting at a small value of $K$ is more important in practice (Chen et al. 2023).
3.2 Empirical Results
In this subsection, we conduct empirical experiments to validate the relationship between Recall@$K$ metric on fully-exposed data and Recall@$\overline{K}$ metric on randomly-exposed data.
We train multiple AutoDebias (Chen et al. 2021) models w.r.t. different hyperparameters and evaluate their Recall@$\overline{K}$ on $D_{rand}$ and Recall@$K$ on $D_{full}$. Then we calculate Pearson and Kendall coefficients between Recall@$\overline{K}$ and Recall@$K$ to measure their linear correlation and ranking correlation (Abdi 2007), respectively. The results are shown in Figure 3. We highlight the value of $K$ when each curve reaches its maximum correlation coefficient as $K_{max}$. From the Figure 3, we have the following observations:
@@ 3.2 Empirical Results
  • When fixing $\overline{K}$, the value of $K_{max}$ decreases as the sample number $\overline{N}$ increases. As shown in Figure 3a and Figure 3c, the inverse relationship between $K_{max}$ and $\overline{N}$ is consistent with the conclusion in Eq. (3) that $K=\frac{N}{\overline{N}}\cdot\overline{K}$. This is intuitively understandable since an increase in the sample size $\overline{N}$ of $D_{rand}$ leads to a stronger and more reliable representation of the fully-exposed dataset $D_{full}$. However, in practice, scaling up the sample size is often impractical due to substantial economic constraints.
@@
  • When fixing $\overline{K}$, the highest Kendall correlation coefficient increase as sample number $\overline{N}$ increase. Figure 3c shows that when sample number $\overline{N}$ is small, Kendall correlation coefficient will be small for any Recall@$K$. Low Kendall correlation coefficient means that if model A has a higher Recall@$\overline{K}$ value on $D_{rand}$ than model B, it does not necessarily indicate that A performs better than B with Recall@$K$ on $D_{full}$. Therefore, it is unreliable to use the Recall@$\overline{K}$ on $D_{rand}$ for model selection.
@@
In conclusion, we prove from both theoretical and experimental aspects that there is a relationship between Recall@$K$ on the fully-exposed dataset and Recall@$\overline{K}$ on the randomly-exposed dataset. Specifically, Recall@$\overline{K}$ is associated with Recall@$K$ when the value of $K$ is relatively large. Given that the gold standard for debiasing recommendation desires Recall@$K$ with small values of $K$ (e.g., $K=50$), using traditional evaluation scheme on randomly-exposed dataset directly is problematic.
4 Entrapment Study
[…]
4.1 Hard Negative Sampling
In this subsection, we begin by introducing the concept of hard negative sampling. Subsequently, we present a theorem delineating the relationship between hard negative sampling and the optimization of Recall@$K$ on $D_{full}$ with the desired value of $K$.
Hard negative sampling refers to oversampling high-scored negative items for training. Taking the representative method DNS($H$, $U$) (Shi et al. 2023) as an example, it first uniformly selects $U$ negative samples and then selects the top $H$ hard samples with the highest prediction scores by the model for training. Particularly, as $H/U$ decreases, it implies that the selected hard negative samples become more challenging for the model. We thus define $H/U$ as the hardness of hard negative sampling.
We can present a theorem elucidating the relationship between hard negative sampling and optimizing Recall@$K$ on $D_{full}$.
Theorem 2.
When the error function $\delta$ is the BCE loss function, the model equipped with hard negative sampling DNS($H$, $U$) approximately optimizes Recall@$K$ metric. Meanwhile, the larger $H/U$ is, the Recall@$K$ metric with a larger $K$ on $D_{full}$ is optimized.
The proof can be found in Appendix A.3. Based on the Theorem 2, we conclude that the hard negative sampling approach could optimize Recall@$K$ with large values of $K$ on $D_{full}$ by using large $H/U$.
Table 2: The statistics of datasets.
Dataset ##\##Users ##\##Items Biased train test
Yahoo!R3 15,400 1,000 311,704 48,600 ($D_{rand}$)
KuaiRec 1411 3327 1,934,404 2,742,166 ($D_{full}$)
Table 3: The comparison between NegS and representative methods with debiasing techniques under the traditional evaluation scheme on $D_{rand}$. The best results are in bold and the second best are underlined. The results are statistically significant with $p<$ 0.01.
recall@$\overline{5}$ MF LightGCN
Yahoo!R3 KuaiRec Yahoo!R3 KuaiRec
Bias 0.7085 0.3546 0.7883 0.4859
IPS 0.6572 0.3102 0.7584 0.4889
DR 0.7195 0.5465 0.7576 0.4849
AutoDebias 0.8085 0.5482 0.8046 0.5181
NegS 0.8330 0.5358 0.8141 0.5288
[…]
4.2.1 Experimental Setup.
To assess the reliability of the traditional evaluation scheme on $D_{rand}$, we use two widely-used datasets for our following experiments: Yahoo!R3 and KuaiRec. The statistics related to these datasets can be found in Table 2. The KuaiRec is the only existing full-exposed dataset that also can be used to construct randomly-exposed set and training set. However, Yahoo!R3 only consists of a biased set for training and an randomly-exposed set for testing. We utilize two representative recommendation models, Matrix Factorization (MF) (Koren, Bell, and Volinsky 2009) and graph neural network model LightGCN (He et al. 2020), as the backbones. We compare NegS with following debiasing methods: IPS (Schnabel et al. 2016), DR (Wang et al. 2019), AutoDebias (Chen et al. 2021). All the models are optimized by Adam optimizer. AutoDebias uses random search to find the best hyperparameters while other models use grid search. The search space of learning rate and weight decay are [1e-4, 1e-3, 1e-2, 1e-1]. For the proposed NegS, we tune $H/U$ and $U$ within the ranges of [0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1] and [50,100,200] respectively. The detialed setup can be found in Appendix A.7.
4.2.2 Observations.
We compare the NegS to debiasing baselines using the traditional evaluation scheme on $D_{rand}$ in Table 3. For NegS, we utilize a large $H/U$ to obtain the best results. As shown in Table 3, we highlight that the NegS exhibits comparable or even better performance compared to the debiasing models on Recall@$\overline{K}$ metric on two datasets. From this example, we can infer that a model that optimizes the Recall@$K$ metric with a large $K$ on $D_{full}$, even without any debiasing technologies, can also achieve remarkable performance under the traditional evaluation scheme on $D_{rand}$. This shows that the traditional evaluation scheme on $D_{rand}$ cannot accurately identify effective debiasing models. As a result, the conclusions drawn from evaluating debiasing techniques in this manner may be unreliable, necessitating significant effort in future work for further investigation.
5 URE scheme
Figure 4: Illustration of the proposed URE scheme on unbiasedly estimating Recall@$K$ on $D_{full}$.

### v2

# Debias Can Be Unreliable: Mitigating Bias in Evaluating Debiasing Recommendation

@@ Abstract.
Recent work has improved recommendation models remarkably by equipping them with debiasing methods. Due to the unavailability of fully-exposed datasets, most existing approaches resort to randomly-exposed datasets as a proxy for evaluating debiased models, employing traditional evaluation scheme to represent the recommendation performance. However, in this study, we reveal that traditional evaluation scheme is not suitable for randomly-exposed datasets, leading to inconsistency between the Recall performance obtained using randomly-exposed datasets and that obtained using fully-exposed datasets. To bridge the gap, we propose the Unbiased Recall Evaluation (URE) scheme, which adjusts the utilization of randomly-exposed datasets to unbiasedly estimate the true Recall performance on fully-exposed datasets. We provide theoretical evidence to demonstrate the rationality of URE and perform extensive experiments on real-world datasets to validate its soundness. Our code and datasets are available at https://anonymous.4open.science/r/URE-6B64/.
Debiasing Recommendation; Evaluation
journalyear: 2025copyright: acmlicensedconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3715548isbn: 979-8-4007-1331-6/2025/04ccs: Information systems Recommender systems
1. Introduction
@@
Recommendation models suffer from various biases leading to deviations from true user preferences even harmful impacts to Web services (Chen et al., 2023; Zhang et al., 2021; Schnabel et al., 2016; Li et al., 2023; Pan et al., 2023). Undoubtedly, user’s true preferences can be accurately measured by the fully-exposed dataset which includes users’ feedback (e.g., ratings) across all items. The gold standard for evaluating debiasing techniques is thus utilizing such a fully-exposed dataset to assess the performance of recommendations (Gao et al., 2022a). However, due to the unavailability of fully-exposed datasets in practice (Gao et al., 2022b), randomly-exposed datasets, which contain users’ feedback on some randomly selected items, are widely employed for debiasing evaluation (Schnabel et al., 2016; Liu et al., 2020; Rosenfeld et al., 2017). Notably, the amount of feedback in randomly-exposed datasets is extremely limited, because randomly exposing items to collect user feedback is highly expensive, and the collection process can severely damage a platform’s revenue.
@@
Previous work calculate widely-used Recall metric on randomly-exposed datasets in the same way as on fully-exposed datasets, which we refer to as the traditional evaluation scheme in subsequent descriptions. Nevertheless, we claim that traditional evaluation scheme is not suitable for computing Recall metric on randomly-exposed datasets when evaluating debiasing recommendation. This is due to the significant disparities between the Recall obtained from the randomly-exposed dataset and those derived from the fully-exposed dataset. To comprehensively investigate the underlying factors contributing to the inconsistency, we undertake a theoretical analysis (cf. Theorem 2) of the evaluation performed on a randomly-exposed dataset. Our analysis demonstrates that the Recall@$\overline{K}$ on the randomly-exposed dataset (e.g., Recall@$\overline{5}$) has a weak correlation with the Recall@$K$ using a small value of $K$ on the fully-exposed dataset (e.g., Recall@50). Although the correlation between Recall@$\overline{K}$ and Recall@$K$ could become stronger as the value of $K$ increases, the performance cutting at a small value of $K$ is more important in practice (Chen et al., 2023). In addition, we conduct extensive experiments on real-world datasets to validate such correlation (cf. Figure 1). Thus, the conclusion derived from the traditional evaluation scheme on the randomly-exposed dataset such as Recall@$\overline{5}$ might be unconvincing as it is inadequate to conclude whether previous debiasing methods have effectively mitigated biases.
@@
Towards this goal, we propose the Unbiased Recall Evaluation (URE) scheme. We set the target as estimating the Recall@$K$ on fully-exposed data, i.e., the ratio of items with positive feedback ranked before position $K$. Considering that a randomly-exposed dataset only includes user feedback on randomly selected items, we can treat it as a representative subset of the fully-exposed dataset, wherein both positive and negative feedback are missing at random. In URE, we sort the model’s prediction scores of all candidate items and utilize the score of $(K+1)$-th item as the threshold to calculate the positive ratio on the randomly-exposed dataset, where the positive ratio is the proportion of items with positive feedback in the randomly-exposed dataset that rank before the threshold. Hence averaging such a positive ratio across all users can be used to unbiasedly estimate Recall@$K$ on a fully-exposed dataset. This is due to the models’ prediction information of all candidate items, which is not considered by traditional evaluation scheme. Through theoretical analysis and extensive experiments, we validate that URE can unbiasedly estimate the Recall@$K$ on fully-exposed data with a desired value of $K$.
The main contributions are summarized as follows:
@@
  • We propose the URE, a novel scheme to unbiasedly estimate the target Recall@$K$ performance with randomly-exposed data.
  • We conduct extensive theoretical and empirical analysis, validating the effectiveness of the URE scheme.
2. Preliminary
Given a fully-exposed dataset (denoted as $D_{full}$), we can determine the true performance of debiasing models by calculating the Recall formula, denoted as denoted as Recall@$K$:
(1) $\begin{split}\text{Recall}@K=\frac{|{k\in Rank_{u}+:k\leq K}|}{|Rank_{u}+|},K\in[1,2,...,L],\end{split}$
where the $Rank_{u}+$ compiles the positions of all items with positive feedback within the personalized ranking list recommended by the models. However, The Recall@$K$ is not supported due to the unavailability of $D_{full}$ (The only existing $D_{full}$ for research is KuaiRec (Gao et al., 2022b)). In this case, researchers employ the same formula to compute Recall on a randomly-exposed dataset (denoted as $D_{rand}$). The resulting Recall is referred to as Recall@$\overline{K}$, where $\overline{K}\in[1,2,...,l],l<<L$. This evaluation approach for debiasing models is termed the ”traditional evaluation scheme on $D_{rand}$”.
@@ 2. Preliminary
In this section, we conduct theoretical analysis and empirical validation to discover why the traditional evaluation scheme on $D_{rand}$ is unreliable by inspecting the relationship between Recall@$K$ on $D_{full}$ and Recall@$\overline{K}$ on $D_{rand}$.
3.1. Theoretical Guarantee
In this subsection, we give the theoretical derivation of the relationship between Recall@$K$ on $D_{full}$ and Recall@$\overline{K}$ on $D_{rand}$.
@@ 3.1. Theoretical Guarantee
(2) $E_{\mathcal{R}}\left[E_{\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}]-\text{Recall@}K\right]=0,$
@@
Proof.
(3) $E_{\mathcal{R}}\left[\text{Recall}@K\right]=\sum_{x=1}^{min(K,N^{+})}\frac{x}{N^{+}}\frac{C^{x}_{N^{+}}C_{N-N^{+}}^{K-x}}{C_{N}^{K}}$
(4) $=\frac{K}{N}\left[\sum_{x=1}^{min(K,N^{+})}\left[C_{N^{+}-1}^{x-1}C_{(N-1)-(N^{+}-1)}^{(K-1)-(x-1)}/C^{K-1}_{N-1}\right]\right]$
(5) $=\frac{K}{N}\left[\ C_{N-1}^{K-1}/C_{N-1}^{K-1}\right]=\frac{K}{N}.$
(6) $E_{\mathcal{R}}\left[E_{\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}]\right]=E_{{N^{+}}^{\prime}}\left[E_{\mathcal{R},\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}|{N^{+}}^{\prime}]\right].$
(7) $E_{{N^{+}}^{\prime}}\left[E_{\mathcal{R},\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}|{N^{+}}^{\prime}]\right]$
(8) $=E_{{N^{+}}^{\prime}}\left[\sum_{x=1}^{min(\overline{K},{N^{+}}^{\prime})}\frac{x}{{N^{+}}^{\prime}}\frac{C^{x}_{{N^{+}}^{\prime}}C^{\overline{K}-x}_{\overline{N}-{N^{+}}^{\prime}}}{C_{\overline{N}}^{\overline{K}}}\right]$
(9) $=E_{{N^{+}}^{\prime}}\left[\frac{\overline{K}}{\overline{N}}\frac{\sum_{x=1}^{min(\overline{K},{N^{+}}^{\prime})}C^{x-1}_{{N^{+}}^{\prime}-1}C^{(\overline{K}-1)-(x-1)}_{\overline{N}-{N^{+}}^{\prime}}}{C_{\overline{N}-1}^{\overline{K}-1}}\right]$
(10) $=E_{{N^{+}}^{\prime}}\left[\frac{\overline{K}}{\overline{N}}\right]=\frac{\overline{K}}{\overline{N}}=\frac{K}{N}=E_{\mathcal{R}}[\text{Recall@}K].$
From Theorem 2, we can observe that the Recall@$\overline{K}$ on $D_{rand}$ will exhibit a strong correlation with the Recall@$K$ on $D_{full}$ if the value of $K$ is around $\frac{{N}}{\overline{N}}\cdot\overline{K}$. It should be noted that $\overline{N}$ is much smaller than $N$ since the $D_{rand}$ is a sparsely sampled subset of $D_{full}$. Therefore, given a specific value of $\overline{K}$, the Recall@$\overline{K}$ on $D_{rand}$ only exhibits a strong correlation with the Recall@$K$ with much larger $K$ values on $D_{full}$. However, given that user typically browse only a limited number of items during an online deployment stage, the performance cutting at a small value of $K$ is more important (Chen et al., 2023).
Figure 1. The correlation coefficients between Recall@$\overline{K}$ on $D_{rand}$ and Recall@$K$ on $D_{full}$. (a) The effect of $\overline{N}$, where $\overline{K}$ is fixed to 5. (b) The effect of $\overline{K}$, where $\overline{N}$ is fixed to 80.
3.2. Empirical Results
In this subsection, we conduct empirical experiments to validate the relationship between Recall@$K$ on $D_{full}$ and Recall@$\overline{K}$ on $D_{rand}$.
We train multiple AutoDebias (Chen et al., 2021) models w.r.t. different hyperparameters and evaluate their Recall@$\overline{K}$ on $D_{rand}$ and Recall@$K$ on $D_{full}$. Then we calculate correlation coefficients between Recall@$\overline{K}$ and Recall@$K$ to measure their correlation. The results are shown in Figure 1. We highlight the value of $K$ when each curve reaches its maximum correlation coefficient as $K_{max}$. From the Figure 1, we have the following observations:
@@ 3.2. Empirical Results
  • When fixing $\overline{K}$, the value of $K_{max}$ decreases as the sample number $\overline{N}$ increases. As shown in Figure 1a, the inverse relationship between $K_{max}$ and $\overline{N}$ is consistent with the conclusion in Eq. (2) that $K=\frac{N}{\overline{N}}\cdot\overline{K}$. This is intuitively understandable since an increase in the sample size $\overline{N}$ of $D_{rand}$ leads to a stronger and more reliable representation of the fully-exposed dataset $D_{full}$. However, in practice, scaling up the sample size is often impractical due to substantial economic constraints.
@@
In conclusion, Recall@$\overline{K}$ is associated with Recall@$K$ when the value of $K$ is relatively large. Given that the gold standard for debiasing recommendation desires Recall@$K$ with small values of $K$ (e.g., $K=50$), Recall@$\overline{K}$ is problematic.
4. URE scheme